# ROLCH: Regularized Online Learning for Conditional Heteroskedasticity

### v1

@@ Abstract
Large-scale streaming data are common in modern machine learning applications and have led to the development of online learning algorithms. Many fields, such as supply chain management, weather and meteorology, energy markets, and finance, have pivoted towards using probabilistic forecasts, which yields the need not only for accurate learning of the expected value but also for learning the conditional heteroskedasticity. Against this backdrop, we present a methodology for online estimation of regularized linear distributional models for conditional heteroskedasticity. The proposed algorithm is based on a combination of recent developments for the online estimation of LASSO models and the well-known GAMLSS framework. We provide a case study on day-ahead electricity price forecasting, in which we show the competitive performance of the adaptive estimation combined with strongly reduced computational effort. Our algorithms are implemented in a computationally efficient python package.
keywords:
GAMLSS, probabilistic forecasting, adaptive learning, distributional regression, volatility modelling, volatility forecasting, weighted regression, iteratively reweighted least squares, coordinate descent, electricity price forecasting, EPF
JEL:
C15; C18; C22; C32; C53; C55; C57; G17; Q47
journal: International Journal of Importance
@@ 1 Introduction
Large-scale streaming data are common in modern applications of machine learning and have led to the development of online learning algorithms (Cesa-Bianchi & Orabona, 2021). For processes driven by a high-dimensional co-variate space, regularized algorithms have been presented by, e.g., Monti et al. (2018); Angelosante et al. (2010); Yang et al. (2023, 2010). In many settings, online statistical algorithms are used to issue forecasts. The advent of probabilistic forecasting in many fields, such as supply chain management, weather and meteorology, energy markets, and finance, yields the need not only for accurate learning of the expected value but also for learning the conditional heteroskedasticity (for a review on distributional regression see, e.g. Klein (2024) and Kneib et al. (2023), for probabilistic forecasting see Petropoulos et al., 2022; Gneiting & Katzfuss, 2014; Ziel et al., 2016; Nowotarski & Weron, 2018; Álvarez et al., 2021; Ziel, 2022). However, online learning approaches for the conditional heteroskedasticity remain sparse in the literature (see e.g. Álvarez et al. 2021 for adaptive heteroskedasticity estimation and Priouret et al. 2005; Dahlhaus & Subba Rao 2007; Werge & Wintenberger 2022; Hendrych & Cipra 2018; Cipra & Hendrych 2018 for online learning for conditional heteroskedasticity) and, to the best of our knowledge, there are no regularized online distributional regression approaches suitable for high-dimensional processes available so far.
@@
In this paper, we provide a regularized online learning algorithm for the conditional distribution parameters of the response variable $\mathcal{Y}$ based on a combination of the online coordinate descent algorithm introduced by Angelosante et al. (2010) and Messner & Pinson (2019) and the generalized additive model for location, scale and shape (GAMLSS) models introduced by Rigby & Stasinopoulos (2005); Stasinopoulos & Rigby (2008); Stasinopoulos et al. (2018). Formally, the GAMLSS framework assumes that $i=1,2,...,n$ independent observations $\mathcal{Y}_{i}$ have the probability density function (PDF)
@@
with (up to) four distribution parameters, each of which can be a (linear) function of explanatory variables. The first two parameters, $\mu_{i}$ and $\sigma_{i}$, commonly characterize the location and scale of the distribution, while $\nu_{i}$ and $\tau_{i}$ are commonly denoted as the shape parameters describing the skewness and kurtosis. Rigby & Stasinopoulos (2005) define the original linear parametric GAMLSS as follows:
$\mathcal{Y}_{i}\sim\mathcal{D}\left(\mu_{i},\sigma_{i},\nu_{i},\tau_{i}\right)\Leftrightarrow\mathcal{D}\left(\theta_{i,1},\theta_{i,2},\theta_{i,3},\theta_{i,4}\right)$ (1)
for $i=0,1,...,n$ independent observations of $\mathcal{Y}_{i}$, where $\mathcal{D}$ is any distribution function with (up to) $p$ distribution parameters $\theta_{i,k}$ for $k=1,...,p$, where $p=4$ commonly [^1] Distributions with more than four parameters can be implemented nevertheless. and let $g_{k}(\cdot)$ be a known, monotonic link function relating a distribution parameter to a predictor $\boldsymbol{\eta}_{k}$ and have
@@
where ${\mathcal{X}}^{k}$ is a known design matrix, $\boldsymbol{\beta}_{k}=(\beta_{k,1},....,\beta_{k,J_{k}})^{\top}$ is a parameter vector of length $J_{k}$ and ${\theta_{k}=(\theta_{0,k},\theta_{1,k},...,\theta_{n,k})}$. The linear parametric GAMLSS model, therefore, allows the modelling of all conditional distribution parameters as linear functions of the explanatory variables in ${\mathcal{X}}_{k}$ and is usually fitted using an iteratively reweighted algorithm on the cross-derivatives of the distribution function $\mathcal{D}$. This setup corresponds to the setting in which LASSO estimation was introduced to GAMLSS by Groll et al. (2019). Further regularized estimation approaches have been proposed by Ziel et al. (2021).
@@
After having received the first $n$ observations, we receive a new pair set of data $Y_{n+1}$ and $X_{k,n+1}$. In the online setting, we are interested in updating the parameters of the model $\boldsymbol{\beta}_{k}$. The proposed algorithm uses the RS algorithm proposed by Rigby & Stasinopoulos (2005) in combination with the online coordinate descent (OCD, see Angelosante et al., 2010; Messner & Pinson, 2019) to update the regression coefficients based on the Gramian matrices. We employ online model selection based on information criteria.
@@
  • 2.
    We provide an open-source, ready-to-use Python implementation of the online GAMLSS model. Our implementation is based on numpy and scipy and employs numba just-in-time compilation for high computational efficiency. The code can be found here https://github.com/simon-hirsch/rolch and the package is available from PyPy (Section 3).
  • 3.
@@
Our work opens multiple avenues for future research. First, our implementation is constrained to linear parametric models. The inclusion of smooth terms and splines seems like a worthwhile extension. On a regular, non-adaptive grid, the inclusion is straightforward by adapting the design matrix ${\mathcal{X}}$. However, automatic, online knot selection for penalized splines is a non-trivial extension. Furthermore, the GAMLSS community has developed multiple extensions for the GAMLSS, e.g. using boosting to enhance the predictive performance or copulas to employ GAMLSS models in multivariate settings. Furthermore, the inclusion of autoregressive and cross-moments effects can provide valuable tools for modelling highly complex processes such as weather, electricity markets or supply chain applications. Lastly, a thorough theoretical analysis of the error bounds of the online model compared to the batch solution should further increase the trust in the presented methods.
@@
The remainder of this paper is structured as follows: Section 2 briefly reviews the literature on online estimation of linear models with a focus on online coordinate descent (Section 2.2) and the RS algorithm for the linear parametric GAMLSS (Section 2.1) and presents the online GAMLSS model. Section 3 presents our open-source implementation. The following section presents our forecasting study setting (Section 4), and Section 5 presents the results. Finally, Section 6 concludes the paper.
2 Methodology
The following section introduces our methodology in detail. We first introduce the notation. The following three subsections present the GAMLSS’ RS algorithm in a batch setting (Section 2.1), give an introduction to online estimation for (regularized) linear models (Section 2.2) and, finally, section 2.3 presents our algorithm ROLCH to enable regularized online learning for conditional heteroskedasticity.
@@ 2 Methodology
Generally, we use a subscript $n$ to indicate up to which observation data is available. Let $\boldsymbol{X}_{n}=(X_{n,1},...,X_{n,J})$ denote the $n$-th vector of $J$ explanatory variables received and ${\mathcal{X}}_{n}=(\boldsymbol{X}_{1},...,\boldsymbol{X}_{n})$ the full data set of $n$ and $J$ explanatory variables. In the online setting, rows $\boldsymbol{X}_{n},\boldsymbol{X}_{n+1},...$ are received subsequently. Similarly, ${\mathcal{Y}}_{n}$ denotes all observations of the dependent variable $(Y_{1},...,Y_{n})$ received by the forecaster so far. In the distributional setting, we have up to $k=1,...,p$ distribution parameters and accordingly up to $p$ data sets ${\mathcal{X}}_{n,k}$ containing the $J_{k}$ covariates for the distribution parameter $k$. We also use a numpy-style notation if we refer to a certain subset of matrices. For arbitrary matrix ${\mathcal{A}}$, ${\mathcal{A}}[n,:]$ denotes accessing the $n$-th row, while ${\mathcal{A}}[:,i]$ denotes accessing the $i$-the column. We start indices generally at 0. We denote changes across different iterations of our algorithm using a superscript $[j]$, for example $\widehat{\beta}_{k}^{[j]}$.
@@ 2.1 Batch GAMLSS
Preliminaries
The following exposition of the RS GAMLSS algorithm largely follows Appendix B of Rigby & Stasinopoulos (2005) and Stasinopoulos & Rigby (2008). Let us introduce some notation first. Generally, ${\mathcal{L}}_{\mathcal{D}}$ and $l_{\mathcal{D}}$ denote the likelihood and log-likelihood function of the distribution ${\mathcal{D}}$ given the data $Y_{i}$, $l(Y\mid\boldsymbol{\theta})$, where $\boldsymbol{\theta}=(\theta_{1},...,\theta_{k})$ for $k$ distribution parameters. Let
@@
$\boldsymbol{z}_{k}=\boldsymbol{\eta}_{k}+\boldsymbol{W}_{kk}^{-1}\boldsymbol{u}_{k}$ (4)
be the adjusted dependent variables. $\boldsymbol{W}_{kk}$ is a diagonal, iterative weight matrix, which can have one of the following forms:
$\boldsymbol{W}_{ks}=-\dfrac{\dfrac{\partial^{2}l}{\partial\theta_{k}}}{\dfrac{\partial\boldsymbol{\eta}_{k}}{\partial\theta_{k}}\left(\dfrac{\partial\boldsymbol{\eta}_{s}}{\partial\theta_{s}}\right)^{\top}}\quad\text{or}\quad\boldsymbol{W}_{ks}=-E\left[\dfrac{\dfrac{\partial^{2}l}{\partial\theta_{k}}}{\dfrac{\partial\boldsymbol{\eta}_{k}}{\partial\theta_{k}}\left(\dfrac{\partial\boldsymbol{\eta}_{s}}{\partial\theta_{s}}\right)^{\top}}\right]\quad\text{or}\quad\boldsymbol{W}_{ks}=-\text{diag}\left\{\dfrac{\dfrac{\partial l}{\partial\theta_{k}}}{\dfrac{\partial\boldsymbol{\eta}_{k}}{\partial\theta_{k}}}\dfrac{\dfrac{\partial l}{\partial\theta_{k}}}{\dfrac{\partial\boldsymbol{\eta}_{s}}{\partial\theta_{s}}}\right\}$ (5)
over $i=1,2,...,n$, i.e. the observed (Newton-Rapson), expected (Fisher-Scoring) or product (quasi Newton-Rapson) score function depending on which kind of algorithm is used in the RS algorithm. We use a Newton-Rapson scoring in our implementation. Hence, the adjusted observation vector $\boldsymbol{z}_{k}$ reads:
$\boldsymbol{z}_{k}=\boldsymbol{\eta}_{k}+\dfrac{\dfrac{\partial l}{\partial\theta_{k}}}{\boldsymbol{W}_{kk}\left(\dfrac{\partial\boldsymbol{\eta}_{k}}{\partial\theta_{k}}\right)}=\boldsymbol{\eta}_{k}+\dfrac{\dfrac{\partial l}{\partial\theta_{k}}}{\left(-\dfrac{\dfrac{\partial\boldsymbol{\eta}_{k}}{\partial\theta_{k}}\left(\dfrac{\partial\boldsymbol{\eta}_{s}}{\partial\theta_{s}}\right)^{\top}}{\dfrac{\partial^{2}l}{\partial\theta_{k}}}\right)\left(\dfrac{\partial\boldsymbol{\eta}_{k}}{\partial\theta_{k}}\right)}$ (6)
@@
where $g^{-1}(\boldsymbol{\eta}_{k})$ is the inverse of the link function (Rigby & Stasinopoulos, 2005; Stasinopoulos & Rigby, 2008; Stasinopoulos et al., 2024).
The Batch RS Algorithm
@@
Recall that $k$ is the parameter index and let $m$ and $r$ be the index for the outer and inner iteration of the RS algorithm. [^2] Note that Rigby & Stasinopoulos (2005) define $r$ as the outer iteration index, $i$ as the inner iteration index and also $i$ as in the observation index. However, overloading the observation index $i$ leads to confusion in the online case, hence we deviate deliberately from their notation here. We denote iterations as superscript, i.e. $\theta^{[m,r]}_{k}$ is the value of $\theta$ for the $k$-th distribution parameter at the outer iteration $m$ and the inner iteration $n$. The RS algorithm consists of two cycles. The outer cycle maximises the penalized likelihood with respect to $\boldsymbol{\beta}_{k}$. The inner cycle is the iterative fitting for each distribution parameter. In each calculation of the algorithm, the most recent updated values of all quantities are used. We omit the back-fitting cycle for smooth terms in this description. The RS algorithm, as outlined in Rigby & Stasinopoulos (2005), is as as follows:
  1. 1.
    (Initialisation:) Initialise the fitted values $\theta_{k}^{[0,0]}$ for $k=1,...,p$. Evaluate the linear predictors ${\boldsymbol{\eta}^{[0,0]}_{k}=g_{k}(\theta_{k}^{[0,0]})}$ for $k=0,...,p$.
  2. 2.
    (Outer Cycle:) Start the outer cycle $m=0,1,...$ until convergence. Iterate over $k=1,....,p$ and
    1. (a)
      Start the inner cycle $n=0,1,...$ until convergence:
      • i.
        Evaluate $\boldsymbol{u}_{k}^{[m,r]}$, $\boldsymbol{W}_{kk}^{[m,r]}$ and $\boldsymbol{z}_{k}^{[m,r]}$
      • ii.

### v2

# Online Distributional Regression

@@ Abstract
Large-scale streaming data are common in modern machine learning applications and have led to the development of online learning algorithms. Many fields, such as supply chain management, weather and meteorology, energy markets, and finance, have pivoted towards using probabilistic forecasts, which yields the need not only for accurate learning of the expected value but also for learning the conditional heteroskedasticity and conditional distribution moments. Against this backdrop, we present a methodology for online estimation of regularized, linear distributional models. The proposed algorithm is based on a combination of recent developments for the online estimation of LASSO models and the well-known GAMLSS framework. We provide a case study on day-ahead electricity price forecasting, in which we show the competitive performance of the incremental estimation combined with strongly reduced computational effort. Our algorithms are implemented in a computationally efficient Python package.
Keywords: online learning, time series, GAMLSS, LASSO, electricity price forecasting (EPF)
@@ 1 Introduction
Large-scale streaming data are common in modern applications of machine learning and have led to the development of online learning algorithms (Cesa-Bianchi and Orabona, 2021). For processes driven by a high-dimensional co-variate space, regularized algorithms have been presented by, e.g., Monti et al. (2018); Angelosante et al. (2010); Yang et al. (2023, 2010). In many settings, online statistical algorithms are used to issue forecasts. The advent of probabilistic forecasting in many fields, such as supply chain management, weather and meteorology, energy markets, and finance, yields the need not only for accurate learning of the expected value but also for learning the conditional heteroskedasticity (for a review on distributional regression see, e.g. Klein (2024) and Kneib et al. (2023), for probabilistic forecasting see Petropoulos et al., 2022; Gneiting and Katzfuss, 2014; Ziel et al., 2016; Nowotarski and Weron, 2018; Álvarez et al., 2021; Ziel, 2022). However, online learning approaches for distributional regression remain sparse in the literature and can be grouped as:
  1. 1.
    Adaptive tracking of the variance respective scale parameter (see e.g. Álvarez et al., 2021; Vilmarest and Wintenberger, 2024),
  2. 2.
    Adaptive estimation of the conditional heteroskedasticity (see e.g Priouret et al., 2005; Dahlhaus and Subba Rao, 2007; Werge and Wintenberger, 2022; Hendrych and Cipra, 2018; Cipra and Hendrych, 2018; Wintenberger, 2024),
  3. 3.
    Adaptive conformal prediction approaches (see e.g. Zaffran et al., 2022; Bhatnagar et al., 2023; Gibbs and Candès, 2024; Dutot et al., 2024; Brusaferri et al., 2024),
and, to the best of our knowledge, there are no regularized online distributional regression approaches suitable for high-dimensional covariate processes available so far.
@@
In this paper, we provide a regularized online learning algorithm for the conditional distribution parameters of the response variable $\mathcal{Y}$ based on a combination of the online coordinate descent algorithm introduced by Angelosante et al. (2010) and Messner and Pinson (2019) and the generalized additive model for location, scale and shape (GAMLSS) models introduced by Rigby and Stasinopoulos (2005); Stasinopoulos and Rigby (2008); Stasinopoulos et al. (2018). Formally, the GAMLSS framework assumes that $i=1,2,...,n$ independent observations $\mathcal{Y}_{i}$ have the probability density function (PDF)
@@
with (up to) four distribution parameters, each of which can be a (linear) function of explanatory variables. The first two parameters, $\mu_{i}$ and $\sigma_{i}$, commonly characterize the location and scale of the distribution, while $\nu_{i}$ and $\tau_{i}$ are commonly denoted as the shape parameters describing the skewness and kurtosis. Rigby and Stasinopoulos (2005) define the original linear parametric GAMLSS as follows:
$Y_{i}\sim\mathcal{D}\left(\mu_{i},\sigma_{i},\nu_{i},\tau_{i}\right)\Leftrightarrow\mathcal{D}\left(\theta_{i,1},\theta_{i,2},\theta_{i,3},\theta_{i,4}\right)$ (1)
for $i=0,1,...,n$ independent observations of $Y_{i}$, where $\mathcal{D}$ is any distribution function with (up to) $p$ distribution parameters $\theta_{i,k}$ for $k=1,...,p$, where $p=4$ commonly [^1] Distributions with more than four parameters can be implemented nevertheless. and let $g_{k}(\cdot)$ be a known, monotonic link function relating a distribution parameter to a predictor $\boldsymbol{\eta}_{k}$ and have
@@
where ${\mathcal{X}}^{k}$ is a known design matrix, $\boldsymbol{\beta}_{k}=(\beta_{k,1},....,\beta_{k,J_{k}})^{\top}$ is a parameter vector of length $J_{k}$ and ${\theta_{k}=(\theta_{0,k},\theta_{1,k},...,\theta_{n,k})}$. The linear parametric GAMLSS model, therefore, allows the modelling of all conditional distribution parameters as linear functions of the explanatory variables in ${\mathcal{X}}_{k}$ and is usually fitted using an iteratively reweighted regression on the cross-derivatives of the distribution function $\mathcal{D}$. This setup corresponds to the setting in which LASSO estimation was introduced to GAMLSS by Groll et al. (2019). Further regularized estimation approaches have been proposed by Ziel et al. (2021).
@@
After having received the first $n$ observations, we receive a new pair set of data $Y_{n+1}$ and $X_{k,n+1}$. In the online setting, we are interested in updating the parameters of the model $\boldsymbol{\beta}_{k}$, without recalling all previous observations. We, therefore, consider an algorithm for a strict online setting in which we can discard all observations after updating the model coefficients. Our proposed algorithm uses the GAMLSS’ RS algorithm proposed by Rigby and Stasinopoulos (2005) and exploits its agnosticism towards the actual estimation method. We, therefore, proceed by combining it with the online coordinate descent (OCD, see Angelosante et al., 2010; Messner and Pinson, 2019) to update the regression coefficients based on the Gramian matrices, which can simply be tracked online. We employ online model selection based on information criteria.
@@
  • We provide an open-source, ready-to-use Python implementation of the online GAMLSS model. Our implementation is based on numpy and scipy and employs numba just-in-time compilation for high computational efficiency. The code can be accessed here https://github.com/simon-hirsch/rolch, and the package is available on PyPy (Section 4).
  • We formally analyse the relationship between the GAMLSS and IRLS models for conditional heteroskedasticity. We employ the theoretical results in a simulation study to shed light on the online GAMLSS’ performance in tracking the true coefficients.
@@
Our work opens multiple avenues for future research. First, our implementation is constrained to linear parametric models. The inclusion of smooth terms and splines seems like a worthwhile extension. On a regular, non-adaptive grid, the inclusion is straightforward by adapting the design matrix ${\mathcal{X}}$. However, automatic, online knot selection for penalized splines is a non-trivial extension. Furthermore, the GAMLSS community has developed multiple extensions for the GAMLSS, e.g. using boosting to enhance the predictive performance or copulas to employ GAMLSS models in multivariate settings. Furthermore, the inclusion of autoregressive and cross-moments effects can provide valuable tools for modelling highly complex processes such as weather, electricity markets or supply chain applications. Lastly, a thorough theoretical analysis of the error bounds of the proposed online approximation compared to the batch setting should further increase the trust in the presented methods.
@@
The remainder of this paper is structured as follows: Section 2 briefly reviews the algorithm for the linear parametric GAMLSS (Section 2.1) and methods for the online estimation of (regularized) linear models (Section 2.2). Section 3 presents our main contribution, the online GAMLSS model. Section 4 presents our open-source implementation. The following Sections 5 and 6 present a simulation study on the properties of the online algorithm and a real-world example for energy markets. Finally, Section 7 discusses our results and concludes the paper.
2 Preliminaries
The following section introduces the preliminaries for the online GAMLSS algorithm. After some notation, the following two subsections present the GAMLSS’ RS algorithm in a batch setting (Section 2.1) and give a brief overview of online estimation for (regularized) linear models (Section 2.2). The Algorithms 1 to 3 are the building blocks for the online distributional regression model, hence we review them here. Moreover, we discuss the relationship between GAMLSS and iteratively reweighted least squares (IRLS) for the estimation of mean-variance models (Francq and Zakoian, 2019; Ziel, 2016; Dette and Wagener, 2013), which we will employ later on to illustrate some limitations of the online algorithm.
@@ 2 Preliminaries
Generally, we use a subscript $n$ to indicate up to which observation data is available. Let $\boldsymbol{X}_{n}=(X_{n,1},...,X_{n,J})$ denote the $n$-th row vector of $J$ explanatory variables received and ${\mathcal{X}}_{n}=(\boldsymbol{X}_{1},...,\boldsymbol{X}_{n})^{\top}$ the full data set of $n$ observations and $J$ explanatory variables. In the online setting, rows $\boldsymbol{X}_{n},\boldsymbol{X}_{n+1},...$ are received subsequently. Similarly, ${\mathcal{Y}}_{n}$ denotes all observations of the dependent variable $(Y_{1},...,Y_{n})$ received by the forecaster so far. In the distributional setting, we have $k=1,...,p$ distribution parameters and accordingly up to $p$ data sets ${\mathcal{X}}_{n,k}$ containing the $J_{k}$ covariates for the distribution parameter $k$. We also use a numpy-style notation if we refer to a certain subset of matrices. For arbitrary matrix ${\mathcal{A}}$, ${\mathcal{A}}[n,:]$ denotes accessing the $n$-th row, while ${\mathcal{A}}[:,i]$ denotes accessing the $i$-the column. We start indices generally at 0. We denote changes across different iterations of our algorithm using a superscript $[j]$, for example $\widehat{\beta}_{k}^{[j]}$.
@@ 2.1 Batch GAMLSS
Algorithm
The following exposition of the RS GAMLSS algorithm largely follows Rigby and Stasinopoulos (2005, Appendix B) and Stasinopoulos and Rigby (2008) and is outlined in Algorithm 1. Let us introduce some notation first. Generally, ${\mathcal{L}}_{\mathcal{D}}$ and $l_{\mathcal{D}}$ denote the likelihood and log-likelihood function of the distribution ${\mathcal{D}}$ given the data $Y_{i}$, $l(Y_{i}\mid\boldsymbol{\theta}_{i})$, where $\boldsymbol{\theta}=(\theta_{1},...,\theta_{k})$ for $k$ distribution parameters. Let
@@
$\boldsymbol{z}_{k}=\boldsymbol{\eta}_{k}+\boldsymbol{W}_{k,k}^{-1}\boldsymbol{u}_{k}$ (4)
be the adjusted dependent variables. $\boldsymbol{W}_{k,k}$ is a diagonal, iterative weight matrix, which can have one of the following forms:
$\boldsymbol{W}_{k,s}=-\dfrac{\dfrac{\partial^{2}l}{\partial\theta_{k}}}{\dfrac{\partial\boldsymbol{\eta}_{k}}{\partial\theta_{k}}\left(\dfrac{\partial\boldsymbol{\eta}_{s}}{\partial\theta_{s}}\right)^{\top}}\quad\text{or}\quad\boldsymbol{W}_{k,s}=-E\left[\dfrac{\dfrac{\partial^{2}l}{\partial\theta_{k}}}{\dfrac{\partial\boldsymbol{\eta}_{k}}{\partial\theta_{k}}\left(\dfrac{\partial\boldsymbol{\eta}_{s}}{\partial\theta_{s}}\right)^{\top}}\right]\quad\text{or}\quad\boldsymbol{W}_{k,s}=-\text{diag}\left\{\dfrac{\dfrac{\partial l}{\partial\theta_{k}}}{\dfrac{\partial\boldsymbol{\eta}_{k}}{\partial\theta_{k}}}\dfrac{\dfrac{\partial l}{\partial\theta_{k}}}{\dfrac{\partial\boldsymbol{\eta}_{s}}{\partial\theta_{s}}}\right\}$ (5)
over $i=1,2,...,n$, i.e. the observed (Newton-Raphson), expected (Fisher-Scoring) or product (quasi Newton-Raphson) score function depending on which kind of algorithm is used in the RS algorithm. We use a Newton-Raphson scoring in our implementation. Hence, the adjusted observation vector $\boldsymbol{z}_{k}$ reads:
$\boldsymbol{z}_{k}=\boldsymbol{\eta}_{k}+\dfrac{\dfrac{\partial l}{\partial\theta_{k}}}{\boldsymbol{W}_{k,k}\left(\dfrac{\partial\boldsymbol{\eta}_{k}}{\partial\theta_{k}}\right)}=\boldsymbol{\eta}_{k}+\dfrac{\dfrac{\partial l}{\partial\theta_{k}}}{\left(-\dfrac{\dfrac{\partial^{2}l}{\partial\theta_{k}}}{\dfrac{\partial\boldsymbol{\eta}_{k}}{\partial\theta_{k}}\left(\dfrac{\partial\boldsymbol{\eta}_{s}}{\partial\theta_{s}}\right)^{\top}}\right)\left(\dfrac{\partial\boldsymbol{\eta}_{k}}{\partial\theta_{k}}\right)}$ (6)
@@
where $g^{-1}(\cdot{})$ is the inverse of the link function (Rigby and Stasinopoulos, 2005; Stasinopoulos and Rigby, 2008; Stasinopoulos et al., 2024). Let $m$ and $r$ be the index for the outer and inner iteration of the RS algorithm. [^2] Note that Rigby and Stasinopoulos (2005) define $r$ as the outer iteration index, $i$ as the inner iteration index and also $i$ as in the observation index. However, overloading the observation index $i$ leads to confusion in the online case, hence we deviate deliberately from their notation here. We denote iterations as superscript, i.e. $\theta^{[m,r]}_{k}$ is the value of $\theta$ for the $k$-th distribution parameter at the outer iteration $m$ and the inner iteration $r$. The algorithm consists of two cycles. The outer cycle maximises the penalized likelihood with respect to $\boldsymbol{\beta}_{k}$. The inner cycle is the iterative fitting for each distribution parameter. In each calculation of the algorithm, the most recent updated values of all quantities are used. We omit the back-fitting cycle for smooth terms in this description, since our work is centered on the linear parametric GAMLSS. Note that the fitting in Line 8 is the core estimation of the conditional distribution parameters and that the RS algorithm is agnostic to which (weighted) statistical learning method employed to regress $\boldsymbol{z}_{k}^{[m,r]}$ against the design matrix ${\mathcal{X}}_{k}$ using the iterative weights $\boldsymbol{W}_{k,k}^{[m,r]}$ (Stasinopoulos et al., 2024, see p. 113). This has facilitated the development of multiple extensions to the batch GAMLSS, such as gamlss.boost (Mayr et al., 2012; Hofner et al., 2014), mixtures with Neural Networks (Rügamer et al., 2024) or regularized approaches as in Groll et al. (2019) and gamlss.lasso (Ziel et al., 2021). We will employ this agnosticism and use online coordinate descent (OCD) to update the distributional parameters online.
Input: ${\mathcal{Y}}_{n},{\mathcal{X}}_{k,n},\text{Distribution }\mathcal{D}$
1 Initialise the fitted values $\hat{\theta}_{k}^{[0,0]}$ for $k=1,...,p$.
2 Evaluate the linear predictors ${\hat{\boldsymbol{\eta}}^{[0,0]}_{k}=g_{k}(\hat{\theta}_{k}^{[0,0]})}$ for $k=0,...,p$.
3 for $m=1,...$ until convergence do
4       forall $k\in 1,...,p$ do
5             Start the inner cycle.
6             for $r=0,1,...$ do
7                   Evaluate $\boldsymbol{u}_{k}^{[m,r]}$, $\boldsymbol{W}_{k,k}^{[m,r]}$ and $\boldsymbol{z}_{k}^{[m,r]}$.